# Instabilities in Convnets for Raw Audio

### v3

@@ II-A Moments of the squared Euclidean norm
Let $x\in\mathbb{R}^{N}$ and $w\in\mathbb{R}^{T}$, $T\leq N$. The circular convolution of $x$ and $w$ satisfies $\|x\ast w\|^{2}=w^{\top}Q_{T}xw$, where the entries of the matrix $Q_{T}x$ are given by $Q_{T}x[n,t]=R_{xx}((t-n)\;\mathrm{mod}\;N)$ for each $0\leq n,t<T$. In particular, all diagonal entries of $Q_{T}x$ are equal to $\|x\|^{2}$.
@@
$\|x\ast w_{\!j}\|^{2}=\operatorname{Tr}\left(w_{\!j}^{\top}Q_{T}xw_{\!j}\right)=\operatorname{Tr}\left(Q_{T}xw_{\!j}w_{\!j}^{\top}\right).$ (5)
@@
$\mathbb{E}\big{[}\|x\ast w_{\!j}\|^{2}\big{]}=\operatorname{Tr}\left(Q_{T}x\mathbb{E}\left[w_{\!j}w_{\!j}^{\top}\right]\right)=\sigma^{2}\operatorname{Tr}\left(Q_{T}x\right).$ (6)
By Lemma II.2, $\operatorname{Tr}(Q_{T}x)=T\|x\|^{2}$, hence $\mathbb{E}[\|x\ast w_{\!j}\|^{2}]=\sigma^{2}T\|x\|^{2}$. For the variance, we recall Theorem 5.2 from [23], which states that if $v\sim\mathcal{N}(\mu,\Sigma)$, then for any matrix $M$
@@
We set $v=w_{\!j}$, $\mu=0$, $\Sigma=\sigma^{2}I$, and $M=Q_{T}x$. We obtain:
$\mathbb{V}\big{[}\|x\ast w_{\!j}\|^{2}\big{]}$ $=2\sigma^{4}\operatorname{Tr}\left(Q_{T}x^{2}\right)$
@@
Let $\lambda$ denote the vector of eigenvalues of $Q_{T}x$. Under the same assumptions as Proposition II.3, and given a deviation $\alpha\geq 0$, the probability of $\|\Phi x\|^{2}$ exceeding $(1+\alpha)\|x\|^{2}$ is bounded from above as
@@
The two propositions above have their own merits. Proposition II.3 as direct consequence of Cantelli’s inequality [24] is straightforward and interpretable in terms of the autocorrelation of $x$. Meanwhile, Proposition II.4, based on Chernoff’s inequality [25], is closer to the empirical percentiles, yet is expressed in terms of the eigenvalues of $Q_{T}x$, for which there is no general formula. The proof is more technical and involves ideas from [26] and [27]. Note that in the particular case of full-length filters ($T=N$), $Q_{T}x$ is a circulant matrix: hence, we interpret these eigenvalues as the energy spectral density of the input signal, i.e., $\lambda=|\hat{x}|^{2}$ where $\hat{x}$ is the discrete Fourier transform of $x$. Detailed proofs for both propositions can be found in the appendix.
@@ III-A From quadratic forms to chi-squared distributions
The inner inequalities $\left(\mathbb{E}\left[A\right]\leq 1\leq\mathbb{E}\left[B\right]\right)$ are a direct consequence of Proposition II.1. Regarding the outer inequalities, we perform an eigenvalue decomposition of $Q_{T}x=U\Lambda U^{\top},$ where the columns of $U$ contain the eigenvectors of $Q_{T}x$ as columns and the diagonal matrix $\Lambda$ contains the spectrum of eigenvalues, $\lambda$. For each filter $w_{\!j}$ with $1\leq j\leq J$, let us use the shorthand $y_{j}=U^{\top}w_{\!j}$. By Lemma II.2 we obtain
@@
Since $Q_{T}x$ is a real symmetric matrix, $U$ is an orthogonal matrix. Thus, $y_{j}$ follows the same distribution as $w_{\!j}$
@@
Lemma II.2 implies $\sum_{k=0}^{T-1}\lambda_{k}=\operatorname{Tr}\left(Q_{T}x\right)=T\|x\|^{2}$. Hence

### v4

@@ II-A Moments of the squared Euclidean norm
Let $x\in\mathbb{R}^{N}$ and $w\in\mathbb{R}^{T}$, $T\leq N$. The circular convolution of $x$ and $w$ satisfies $\|x\ast w\|^{2}=w^{\top}Q_{T}(x)w$, where the entries of the matrix $Q_{T}(x)$ are given by $Q_{T}(x)[n,t]=R_{xx}((t-n)\;\mathrm{mod}\;N)$ for each $0\leq n,t<T$. In particular, all diagonal entries of $Q_{T}(x)$ are equal to $\|x\|^{2}$.
@@
$\|x\ast w_{\!j}\|^{2}=\operatorname{Tr}\left(w_{\!j}^{\top}Q_{T}(x)w_{\!j}\right)=\operatorname{Tr}\left(Q_{T}(x)w_{\!j}w_{\!j}^{\top}\right).$ (5)
@@
$\mathbb{E}\big{[}\|x\ast w_{\!j}\|^{2}\big{]}=\operatorname{Tr}\left(Q_{T}(x)\mathbb{E}\left[w_{\!j}w_{\!j}^{\top}\right]\right)=\sigma^{2}\operatorname{Tr}\left(Q_{T}(x)\right).$ (6)
By Lemma II.2, $\operatorname{Tr}(Q_{T}(x))=T\|x\|^{2}$, hence $\mathbb{E}[\|x\ast w_{\!j}\|^{2}]=\sigma^{2}T\|x\|^{2}$. For the variance, we recall Theorem 5.2 from [23], which states that if $v\sim\mathcal{N}(\mu,\Sigma)$, then for any matrix $M$
@@
We set $v=w_{\!j}$, $\mu=0$, $\Sigma=\sigma^{2}I$, and $M=Q_{T}(x)$. We obtain:
$\mathbb{V}\big{[}\|x\ast w_{\!j}\|^{2}\big{]}$ $=2\sigma^{4}\operatorname{Tr}\left(Q_{T}(x)^{2}\right)$
@@
Let $\lambda$ denote the vector of eigenvalues of $Q_{T}(x)$. Under the same assumptions as Proposition II.3, and given a deviation $\alpha\geq 0$, the probability of $\|\Phi x\|^{2}$ exceeding $(1+\alpha)\|x\|^{2}$ is bounded from above as
@@
The two propositions above have their own merits. Proposition II.3 as direct consequence of Cantelli’s inequality [24] is straightforward and interpretable in terms of the autocorrelation of $x$. Meanwhile, Proposition II.4, based on Chernoff’s inequality [25], is closer to the empirical percentiles, yet is expressed in terms of the eigenvalues of $Q_{T}(x)$, for which there is no general formula. The proof is more technical and involves ideas from [26] and [27]. Note that in the particular case of full-length filters ($T=N$), $Q_{T}(x)$ is a circulant matrix: hence, we interpret these eigenvalues as the energy spectral density of the input signal, i.e., $\lambda=|\hat{x}|^{2}$ where $\hat{x}$ is the discrete Fourier transform of $x$. Detailed proofs for both propositions can be found in the appendix.
@@ III-A From quadratic forms to chi-squared distributions
The inner inequalities $\left(\mathbb{E}\left[A\right]\leq 1\leq\mathbb{E}\left[B\right]\right)$ are a direct consequence of Proposition II.1. Regarding the outer inequalities, we perform an eigenvalue decomposition of $Q_{T}(x)=U\Lambda U^{\top},$ where the columns of $U$ contain the eigenvectors of $Q_{T}(x)$ as columns and the diagonal matrix $\Lambda$ contains the spectrum of eigenvalues, $\lambda$. For each filter $w_{\!j}$ with $1\leq j\leq J$, let us use the shorthand $y_{j}=U^{\top}w_{\!j}$. By Lemma II.2 we obtain
@@
Since $Q_{T}(x)$ is a real symmetric matrix, $U$ is an orthogonal matrix. Thus, $y_{j}$ follows the same distribution as $w_{\!j}$
@@
Lemma II.2 implies $\sum_{k=0}^{T-1}\lambda_{k}=\operatorname{Tr}\left(Q_{T}(x)\right)=T\|x\|^{2}$. Hence